A random sample of size 36 was taken from the population with a mean of 3.4 and a variance of 4. Approximately the probability that the sample mean is greater than 3.0 but less than 3.6


1
Expert's answer
2022-04-26T11:55:06-0400

Let "X=" the sample mean, "X\sim N(\mu, \sigma^2/n)."


"P(3.0<X<3.6)=P(X<3.6)-P(X\le 3.0)"

"=P(Z<\dfrac{3.6-3.4}{\sqrt{4}/\sqrt{36}})-P(Z\le\dfrac{3.0-3.4}{\sqrt{4}/\sqrt{36}})"

"=P(Z<0.6)-P(Z\le-1.2)"

"\approx0.7257469-0.1150697"

"\approx 0.610677"